OpenStudy (anonymous):

solve by the substitute method 6x+8y=4 -9x+y=59

OpenStudy (stormfire1):

Original equations: 6x + 8y = 4 -9x+y=59 Isolate y y=59+9x Plug in y in place of x in one of the equations and solve for x 6x+8(59+9x)=4 6x+472+72x=4 78x=-468 x=-6 Plug -6 back into one of the other equations and solve for y. You should end up with y = 5 and x = -6